# Deeper quantum circuits via pseudo-twirling coherent errors mitigation in non-Clifford gates

### v1

@@ Abstract
The conventional circuit paradigm, utilizing a limited number of gates to construct arbitrary quantum circuits, is hindered by significant noise overhead. For instance, the standard gate paradigm employs two CNOT gates for the partial ZZ rotation in the quantum Fourier transform, even when the rotation angle is very small. In contrast, certain devices, such as IBM cross-resonance-based devices, can directly implement such operations using their native interaction, resulting in considerably shorter and less noisy implementations for small rotation angles. Unfortunately, beyond noise (incoherent errors), coherent errors stemming from qubit crosstalk and calibration imperfections render these implementations impractical. In Clifford gates such as the CNOT, these errors can be addressed through Pauli twirling (also known as randomized compiling). However, this technique is not applicable to the short non-Clifford native implementations described above. The present work introduces, analyzes, and experimentally demonstrates a technique called Pseudo Twirling to address coherent errors in general gates and circuits. Additionally, we experimentally showcase that integrating pseudo twirling with a quantum error mitigation method called adaptive KIK enables the simultaneous mitigation of both noise and coherent errors in non-Clifford gates. This advancement paves the way for error mitigation in larger circuits than ever before.
@@
In recent years, quantum error mitigation protocols [1, 2, 3, 4, 5, 6, 7, 8, 9] have significantly boosted the performance of quantum computers [10, 11, 5, 12, 13, 14, 15, 16, 17, 18]. These methods are primarily designed for managing incoherent errors (noise) arising from interactions with the environment. However, in addition to incoherent errors, there are also coherent errors that cannot be addressed using the same methods in a scalable manner. Coherent errors pose a significant concern in the development of quantum computers in the foreseeable future. They are often attributed to calibration errors, drift in calibration parameters, or coherent crosstalk interactions between qubits. While hardware continues to advance, the demand for minimizing coherent errors becomes increasingly crucial as circuits grow in size and complexity.
@@
Currently, the most effective approach to address these errors is a method known as Randomized Compiling (RC) [19, 14, 20, 21], also referred to as Pauli twirling. However, this technique is limited to Clifford gates. For Clifford gates such as the CNOT, iswap, and others, RC involves executing an ensemble of circuits that are equivalent in the absence of coherent errors, but each circuit alters the coherent error in a distinct way. After running these circuits and performing post-processing averaging, the coherent error is mitigated, but the resulting transformation is no longer a unitary. In other words, the coherent error is converted into an incoherent error. This conversion is beneficial as incoherent errors can be addressed with QEM (quantum error mitigation) protocols. As demonstrated experimentally in [22] the integration of RC with a QEM method called ‘adaptive KIK’, was imperative for removing coherent errors and obtain accurate results.
@@
Another interesting feature of the RC method is that it converts a general incoherent error channel into a Pauli channel that has a much simpler structure [19]. Due to the reduced degrees of freedom in the Pauli channel compared to general noise, it becomes feasible to employ additional sparsity techniques and efficiently characterize the noise channel of multi-qubit Clifford gates [17, 23, 20]. This key feature is what facilitates two important QEM techniques, PEC [4] and PEA [10]. PEA presently hold the record of the largest circuit that QEM has been applied to – 127 qubit Ising lattice time evolution experiment [10]. The largest depth probed in this experiment was 60 cnots.
@@
Regrettably, RC is restricted to Clifford gates, rendering PEC and PEA inapplicable to non-Clifford multi-qubit gates. Furthermore, circuits consisting solely of Clifford gates can be efficiently simulated, making it pointless to build a powerful quantum computer just for their execution. Although multi-qubit non-Clifford gates are not part of the conventional universal gate set paradigm, they can prove crucial in specific applications, particularly in the era of noisy quantum computers where circuit depth is limited.
As a first example we bring the $n$-qubit quantum Fourier transform [24] that comprises of single-qubit gates and two-qubit gates that execute ZZ rotations with an angle of $\theta=\pi/2^{n}$ (and multiples of it). As $n$ increases, the gate performs a transformation that is very close to the identity transformation. Such gates can be implemented very rapidly when the correct control signals are available, making them less susceptible to noise. In fact, they can be easily implemented using the native two-qubit cross-resonance interaction [25, 26] exploited in IBM gates, as depicted in Fig. 1a. In contrast, in the gate paradigm, the length of the ZZ rotation gate is determined by the length of two CNOT gates, even when the rotation angle is minute (see 1b).
The second example pertains to the simulation of physical systems. For instance, the transverse Ising lattice model [27] also necessitates fractional ZZ gates. Although the implementation of a native fractional ZZ non Clifford gate is not an obstacle (e.g., in the IBM platform), its incompatibility with RC poses a significant problem. The coherent errors in this gate currently cannot be effectively addressed with existing methods.
@@
Figure 1: Non-Clifford vs. Clifford implementation of quantum gates. (a) The useful two-qubit gate $U_{zz}(\theta)=e^{-i\theta Z_{1}Z_{2}}$ can be implemented compactly in the IBM platform using a non-Clifford gate $e^{-i\theta X_{1}Z_{2}}$, directly generated from the native interaction between the qubits. Conversely, the current Clifford-based implementation (b) is substantially longer, with each box representing a CNOT gate. Consequently, this implementation exhibits considerably higher levels of noise. However, currently, there is no technique for addressing coherent errors in non-Clifford gates, providing an advantage to the more noisy implementation depicted in (b). Our objective is to develop a protocol for mitigating coherent errors in compact, low-noise non-Clifford gates, as shown in (a).
@@
In this work we present a new twirling scheme called ‘pseudo twirling’ (PST). This approach offers three noteworthy characteristics: 1) PST is universally applicable to all gates, including non-Clifford gates. 2) In contrast to RC (Pauli twirling), PST distinguishes between two types of coherent errors: over(under)-rotation errors and uncontrolled coherent errors. Over-rotations typically arise from imperfect initial calibration or a calibration drift. The uncontrolled errors stem from various mechanism ranging from crosstalk between qubits to unwanted interaction generated by the driving field. PST selectively transforms uncontrolled errors into noise, leaving calibration errors unaffected. 3) PST does not produce a Pauli error channel; instead, it appears to generate a distinct, beneficial noise channel. In this paper we present and study these three features. By integrating PST with QEM this research can have an immediate impact on the capabilities of quantum computers.
@@
With the restriction of applying twirling exclusively to Clifford gates removed, PST can be employed within individual components of a gate (intra-gate PST). As we show later, this facilitates more robust mitigation of coherent errors since coherent errors are addressed before they have a chance to accumulate. Additionally, PST’s flexibility allows for its application across an entire circuit, where the twirling occurs at the edges. Although less effective than intra-gate PST, it can capture coherent errors occurring on a scale exceeding that of multi-qubit gates.
We would like to highlight that, after the completion of the theory presented in this paper, we became aware that IBM is also exploring a similar concept. However, we lack specific details regarding their idea and how it was exploited. Our work was conducted entirely independently, and we did not receive any technical or scientific input from IBM or any other entity.
@@
For convenience we shall use Liouville space representation where the density matrix is flattened in column vector $\left|\rho\right\rangle$ and the evolution is described by superoperators operating on this vector from the left $\left|\rho(t)\right\rangle=K(t)\left|\rho(0)\right\rangle$. See Appendix I for more details.
@@
Pauli matrices in Liouville space - Let us denote by $P_{\alpha}$ the product of single-qubit Pauli matrices $\sigma_{i}$ such that $P_{\alpha}\in\{\sigma_{k}\otimes\sigma_{l}\otimes\sigma_{m}...\}_{k,l,m..\in\{0,x,y,z\}}^{n}$. In Hilbert space the Pauli evolution operator and Pauli Hamiltonian are the same. In Liouville space the two expression differ. A Pauli Hamiltonian in Liouville space is given by $\mathcal{H}_{P_{\alpha}}=P_{\alpha}\otimes I-I\otimes P_{\alpha}^{t}$ while a Pauli evolution operator has the form $P_{\alpha}\otimes P_{\alpha}^{*}$ in Liouville space (see Appendix I). The two are related via the relation $\mathcal{P}_{\alpha}=exp(-i\frac{\pi}{2}\mathcal{H}_{P_{\alpha}})$. While some of the derivations in this work can be carried our in Hilbert space for clarity and consistency we carry out all calculations in Liouville space. A single exception will be discussed when introduced.
I Description of the PST Methodology
I.1 Regular Randomized Compiling
@@
For clarity, we start by describing the regular randomized compiling/Pauli twirling scheme which is applicable only for Clifford gates. Let $U_{\text{cliff}}$ stand for the evolution operator of an ideal Clifford gate and let $K_{\text{cliff}}$ stand for the evolution operator of the noisy realization of $U_{\text{cliff}}$. $K_{\text{cliff}}$ may encompass both noise and coherent errors, but, for the moment, we will concentrate on the noise-free scenario where only coherent errors take place.
@@
As shown in Fig. 2a, in RC, $K_{\text{cliff}}$ is replaced by an averaged operator, with each element "twirled" by an n-qubit Pauli operators $\mathcal{P}_{\alpha}$ and $\mathcal{P}^{\prime}_{\alpha}$.
$K_{\text{cliff}}^{RC}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}^{\prime}K_{\text{cliff}}\mathcal{P}_{\alpha}.$
@@
$\mathcal{P}^{\prime}_{\alpha}$ is chosen in such a way that it yields $K_{\text{cliff}}^{RC}=U_{\text{cliff}}$ in the absence of coherent or incoherent errors. This condition of invariance can be expressed as $\mathcal{P}_{\alpha}^{\prime}U_{\text{cliff}}\mathcal{P}_{\alpha}=U_{\text{cliff}}$, leading to the determination of $\mathcal{P}_{\alpha}^{\prime}$ as $\mathcal{P}_{\alpha}^{\prime}=U_{\text{cliff}}\mathcal{P}_{\alpha}U_{\text{cliff}}^{\dagger}$. Therefore, $P_{\alpha}^{\prime}$ is defined based on the selected $\mathcal{P}_{\alpha}$. Up to his point it appears as we have not used the fact that $U_{\text{cliff}}$ is Clifford. However, since Pauli operators are stabilizers of the Clifford gates, it is guaranteed that if $\mathcal{P}_{\alpha}$ is a Pauli operator, then $\mathcal{P}_{\alpha}^{\prime}$ is also a Pauli operator. Conversely, if the ideal gate is non-Clifford, $\mathcal{P}_{\alpha}^{\prime}$ may involve multi-qubit interactions, introducing noise and coherent errors comparable to those RC aims to correct. These additional errors could potentially worsen the twirled operator. Thus, a significant advantage of Pauli Twirling lies in the experimental simplicity of the twirling operations, which involve very low levels of coherent and incoherent errors.
@@
Notably, the full number of twirlings for $n$-qubit gate, $2^{2n}$, can swiftly grow to a very substantial value. This becomes particularly evident when considering sequences of twirled gates. Fortunately, there is no need to sum over all possible Pauli operators. It suffices to sample a subset from the set $\{\mathcal{P}_{\alpha}\}{}_{\alpha=1}^{2^{2n}}$ to obtain the average value. Nevertheless, for the sake of simplicity we will maintain the use of $2^{n}$.
@@
To gain a better grasp of how RC operates, let us express $K_{\text{cliff}}$ as $K_{\text{cliff}}=U_{\text{cliff}}N$, where $N$ represents the error channel that encompasses all errors. This decomposition always exists when the dynamics of the gate constitute a completely positive map. Substituting it into the twirling formula, we obtain:
$K_{\text{cliff}}^{RC}$ $=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}^{\prime}U_{\text{cliff}}N\mathcal{P}_{\alpha}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}^{\prime}U_{\text{cliff}}\mathcal{P}_{\alpha}\mathcal{P}_{\alpha}N\mathcal{P}_{\alpha}$
$=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}U_{\text{cliff}}\mathcal{P}_{\alpha}N\mathcal{P}_{\alpha}=U_{\text{cliff}}[\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}N\mathcal{P}_{\alpha}].$
@@
Hence, twirling a Clifford gate with Pauli operators is equivalent to twirling its noise channel. Note however that for the noise twirling $\mathcal{P}^{\prime}_{\alpha}=\mathcal{P}{}_{\alpha}$. Furthermore, we can utilize the fact that Pauli operators form a complete basis, allowing any matrix $A$ to be represented as $A=\sum_{\alpha,\beta}A_{\alpha,\beta}P_{\alpha}\otimes P_{\beta}^{t}$ where ${}^{\prime}t^{\prime}$ is the transpose operation. It can be shown that, after Pauli twirling, $A$ transforms into
$A^{RC}=[\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}P_{\alpha}AP_{\alpha}]=\sum_{\alpha}A_{\alpha,\alpha}P_{\alpha}\otimes P_{\alpha}^{t}.$
@@
In this transformation, only the diagonal elements in the diagonal basis remain. This resulting channel is referred to as a Pauli error channel, and it corresponds to randomly applying a Pauli unitary $P_{\alpha}$ with a probability of $N_{\alpha,\alpha}$. Thus, regardless of the original structure of $N$, Pauli twirling turns it into a Pauli channel.
@@
Figure 2: (a) Randomized compiling (RC, Pauli twirling) replaces the original CNOT with a sum of Pauli-twirled CNOT gates. RC can only mitigate coherent errors in Clifford gates. (b) Our proposed Pseudo Twirling (PST) scheme can mitigate uncontrolled coherent errors, such as crosstalk, in a general gate. This is achieved by modifying the signs of some of the pulses in the driving field that generates the gate according to the rule $H_{d}(t)\to P_{\alpha}H_{d}(t)P_{\alpha}$. (c) Because PST is not restricted to Clifford gates, it can be implemented at various levels of a circuit: at the entire circuit level (orange lines), gate-level (green lines), and intra-gate level (purple lines). One of our objectives is to study various PST protocols and explore their optimal application in quantum circuits.
I.2 Introducing the Pseudo Twirling Protocol

### v2

# Pseudo Twirling Mitigation of Coherent Errors in non-Clifford Gates

@@ Abstract
The conventional circuit paradigm, utilizing a limited number of gates to construct arbitrary quantum circuits, is hindered by significant noise overhead. For instance, the standard gate paradigm employs two CNOT gates for the partial CPhase rotation in the quantum Fourier transform, even when the rotation angle is very small. In contrast, some quantum computer platforms can directly implement such operations using their native interaction, resulting in considerably shorter and less noisy implementations for small rotation angles. Unfortunately, coherent errors stemming from qubit crosstalk and calibration imperfections render these implementations impractical. In Clifford gates such as the CNOT, these errors can be addressed through Pauli twirling (also known as randomized compiling). However, this technique is not applicable to the non-Clifford native implementations described above. The present work introduces, analyzes, and experimentally demonstrates a technique called ‘Pseudo Twirling’ to address coherent errors in general gates and circuits. Additionally, we experimentally showcase that integrating pseudo twirling with a quantum error mitigation method called ‘Adaptive KIK’ enables the simultaneous mitigation of both noise and coherent errors in non-Clifford gates. Due to its unique features pseudo twirling could become a valuable asset in enhancing the capabilities of both present and future NISQ devices.
I Introduction
@@
In recent years, quantum error mitigation protocols (QEM) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] have significantly boosted the performance of quantum computers [11, 12, 5, 13, 14, 15, 16, 17, 18, 19]. These methods are primarily designed for managing incoherent errors (noise) arising from interactions with the environment. However, in addition to incoherent errors, there are also coherent errors that cannot be addressed using the same methods in a scalable manner. Coherent errors pose a significant concern in the development of quantum computers in the foreseeable future. They are often attributed to calibration errors, drift in calibration parameters, or coherent crosstalk interactions between qubits. While hardware continues to advance, the demand for minimizing coherent errors becomes increasingly crucial as circuits grow in size and complexity.
@@
Currently, the most effective approach to address these errors is a method known as Randomized Compiling (RC) [20, 15, 21, 22, 23], also referred to as Pauli twirling. However, this technique is limited to Clifford gates. For Clifford gates such as the CNOT, iSWAP, and others, RC involves executing an ensemble of circuits that are equivalent in the absence of coherent errors, but each circuit alters the coherent error in a distinct way. After running these circuits and performing post-processing averaging, the coherent error is mitigated, but the resulting transformation is no longer a unitary. In other words, the coherent error is converted into an effective incoherent error. This conversion is beneficial as incoherent errors can be addressed with QEM protocols. As demonstrated experimentally in [24] the integration of RC with a QEM method called ‘Adaptive KIK’, was imperative for removing coherent errors and obtain accurate results.
@@
Another interesting feature of RC is that it converts a general incoherent error channel into a Pauli channel that has a much simpler structure [20]. Due to the reduced degrees of freedom in the Pauli channel compared to general noise, it becomes feasible to employ additional sparsity techniques and efficiently characterize the noise channel of multi-qubit Clifford gates [18, 25, 21]. This key feature facilitates two important QEM techniques, PEC [4] and PEA [11]. PEA presently holds the record of the largest circuit that QEM has been applied to – 127 qubit Ising lattice time-evolution experiment [11]. The largest depth probed in this experiment was sixty cnots.
Unfortunately, RC is restricted to Clifford gates, rendering PEC and PEA inapplicable to non-Clifford multi-qubit gates. A recent work [26] extends these methods to non-Clifford gates. In contrast to the Clifford case, the sampling overhead per gate can be substantial regardless of how small the coherent error is. In [27] a partial RC was applied by using only the Paulis that commutes with the non-Clifford gate. This partial twirling can only address coherent errors that happen to anti-commute with the selected Pauli subset. Thus, it is not possible to rely on this method if the coherent error is of unknown type, or if it known to commute with the selected Pauli subset.
@@
Circuits consisting solely of Clifford gates can be efficiently simulated, making it pointless to build a powerful quantum computer just for their execution. Although multi-qubit non-Clifford gates are not part of the conventional universal gate set paradigm, they can prove crucial in specific applications, particularly in the era of noisy quantum computers where circuit depth is a limiting factor.
As a first example, we bring the time-evolution of the transverse Ising lattice model [28]. The circuits implementing this evolution require $R_{zz}(\theta)$ gates that can be directly implemented very rapidly, leading to significantly lower noise levels. For example, in [27] the IBM cross-resonance interaction [29, 30, 31], was facilitated to directly implement $R_{zz}(\theta)$ using two additional single qubit gates as depicted in Fig. 1a. In contrast, in the cNOT gate paradigm, the length of the ZZ rotation gate is determined by the length of two CNOT gates, even when the rotation angle is minute (see 1b). The direct implementation in [27] has enabled the execution of Ising time evolution that are four times longer compared to the double-CNOT implementation (Fig. 1b). In trapped ions, a similar statement holds for the Mølmer–Sørensen gate [32]. Although in this case the direct non-Clifford gate may not be shorter than a single CNOT gate, it is at least twice shorter than the double-CNOT implementation.
The second example pertains to the $n$-qubit quantum Fourier transform [33] that comprises of single-qubit gates and two-qubit gates that execute CPhase($\theta$) rotations with an angle of $\theta=\pi/2^{n}$ (and multiples of it). As $n$ increases, the gate performs a transformation that is very close to the identity transformation. As such, a substantial advantage in terms of noise is expected when using the direct implementation (the CPhase($\theta$) analogue of Fig. 1a) instead of the double-CNOT implementation. Hence, a method for mitigating coherent errors in non-Clifford gates would benefit the implementation quantum Fourier transforms and other circuits as well.
@@
Figure 1: Non-Clifford vs. Clifford implementation of quantum gates. (a) The useful two-qubit gate $R_{zz}(\theta)=e^{-i\frac{1}{2}\theta Z_{1}Z_{2}}$ can be implemented compactly in the IBM platform using a non-Clifford gate $R_{zx}(\theta)=e^{-i\frac{1}{2}\theta Z_{1}X_{2}}$, generated by the cross-resonance interaction. Conversely, the current Clifford-based implementation (b) is substantially longer, with each box representing a CNOT gate. Consequently, this implementation exhibits considerably higher levels of noise. However, currently, there is no technique for addressing coherent errors in non-Clifford gates, providing an advantage to the more noisy implementation depicted in (b). In this work we present a protocol for mitigating coherent errors in compact, low-noise non-Clifford gates, as shown in (a).
@@
In this work we present a new twirling scheme called ‘Pseudo Twirling’ (PST). This approach offers three noteworthy characteristics: 1) PST is equally applicable to both Clifford and non-Clifford gates. 2) In contrast to RC (Pauli twirling), PST distinguishes between controlled over(under)-rotation errors and all other controlled or uncontrolled coherent errors. Over-rotations typically arise from imperfect initial calibration or a calibration drift. Others errors stem from various mechanism ranging from crosstalk between qubits to parasitic higher-order interactions created the driving fields. PST transforms all coherent error excluding controlled rotation errors - a fact that can be exploited for high accuracy calibration. 3) PST does not produce a Pauli error channel; instead, it simplifies the noise in a different manner. In this paper we present and study these three features. The integration of PST with QEM, as demonstrated in the present work, can have an immediate impact on the capabilities of quantum computers.
@@
With the restriction of applying twirling exclusively to Clifford gates removed, PST can be employed within individual components of a gate (intra-gate PST). As we show later, this facilitates more effective mitigation of coherent errors since coherent errors are addressed before they have a chance to accumulate. Additionally, PST’s flexibility allows for its application across an entire circuit, where the twirling occurs at the edges. Although this ‘Edges PST’ is less effective than intra-gate PST, it can capture coherent errors occurring on a scale exceeding that of multi-qubit gates.
@@
For convenience we shall use Liouville space representation where the density matrix is flattened into a column vector $\left|\rho\right\rangle$ and the evolution is described by superoperators operating on this vector from the left $\left|\rho(t)\right\rangle=\mathcal{K}(t)\left|\rho(0)\right\rangle$. If the dynamics is unitary the evolution of the density matrix $\left|\rho(t)\right\rangle$ is determined by Schrödinger-like equation where
$\frac{d}{dt}\left|\rho(t)\right\rangle$ $=-i\mathcal{H}(t)\left|\rho(t)\right\rangle,$ (1)
$\mathcal{H}(t)$ $=H(t)\otimes I-I\otimes H(t)^{\intercal}.$ (2)
[…]
$\mathcal{U}(t)=U(t)\otimes U(t)^{*},$ (3)
@@
where $U(t)$ is the evolution operator in Hilbert space. Pauli matrices in Liouville space can either appear as Hamiltonians or as unitaries, but unlike in Hilbert space, these two forms differ. Let us denote by $P_{\alpha}$ the tensor product of single-qubit Pauli matrices $\sigma_{i}$ such that $P_{\alpha}\in\{\sigma_{k}\otimes\sigma_{l}\otimes\sigma_{m}...\}_{k,l,m..\in\{0,x,y,z\}}$. A Pauli Hamiltonian in Liouville space is given by $\mathcal{H}_{\alpha}=P_{\alpha}\otimes I-I\otimes P_{\alpha}^{\intercal}$, while according to Eq. (3), a Pauli evolution operator has the form $\mathcal{P}_{\alpha}=P_{\alpha}\otimes P_{\alpha}^{*}$ in Liouville space. These two are related via $\mathcal{P}_{\alpha}=exp(-i\frac{\pi}{2}\mathcal{H}_{\alpha})$. See Appendix I for more details.
II Description of the PST methodology
II.1 Regular Randomized Compiling
@@
For clarity, we start by describing the regular randomized compiling/Pauli twirling scheme, which is applicable only for Clifford gates. Let $\mathcal{U}_{\text{cliff}}$ stand for the evolution operator of an ideal Clifford gate, and let $\mathcal{K}_{\text{cliff}}$ stand for the evolution operator of the noisy realization of $\mathcal{U}_{\text{cliff}}$. $\mathcal{K}_{\text{cliff}}$ may encompass both noise and coherent errors, but, for the moment, we will concentrate on the noise-free scenario where only coherent errors take place.
@@
As shown in Fig. 2a, in RC, $\mathcal{K}_{\text{cliff}}$ is replaced by an averaged operator, with each element “twirled” by an n-qubit Pauli operators $\mathcal{P}_{\alpha}$ and $\mathcal{P}^{\prime}_{\alpha}$
$\mathcal{K}_{\text{cliff}}^{RC}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}^{\prime}\mathcal{K}_{\text{cliff}}\mathcal{P}_{\alpha}.$ (4)
@@
$\mathcal{P}^{\prime}_{\alpha}$ is chosen in such a way that it yields $\mathcal{K}_{\text{cliff}}^{RC}=\mathcal{U}_{\text{cliff}}$ in the absence of coherent and incoherent errors. This condition of invariance can be expressed as $\mathcal{P}_{\alpha}^{\prime}\mathcal{U}_{\text{cliff}}\mathcal{P}_{\alpha}=\mathcal{U}_{\text{cliff}}$, leading to the determination of $\mathcal{P}_{\alpha}^{\prime}$ as $\mathcal{P}_{\alpha}^{\prime}=\mathcal{U}_{\text{cliff}}\mathcal{P}_{\alpha}\mathcal{U}_{\text{cliff}}^{\dagger}$. Therefore, $\mathcal{P}_{\alpha}^{\prime}$ is determined by the selected $\mathcal{P}_{\alpha}$. Up to his point it appears as if we have not used the fact that $\mathcal{U}_{\text{cliff}}$ is a Clifford gate. However, since Clifford gates normalize the Pauli operators, it is guaranteed that if $\mathcal{P}_{\alpha}$ is a Pauli operator, then $\mathcal{P}_{\alpha}^{\prime}$ is also a Pauli operator. Conversely, if the ideal gate is non-Clifford, $\mathcal{P}_{\alpha}^{\prime}$ may involve multi-qubit interactions, introducing noise and coherent errors comparable to those RC aims to correct. These additional errors could potentially degrade the performance of the gate with respect to the non-twirled gate. Thus, a significant advantage of Pauli Twirling lies in the experimental simplicity of the twirling operations, which involve very low levels of coherent and incoherent errors.
@@
Notably, the total number of twirlings for $n$-qubit gate, $2^{2n}$, can swiftly grow to a very substantial value. This becomes particularly evident when considering sequences of twirled gates. Fortunately, there is no need to sum over all possible Pauli operators. It suffices to sample from the set $\{\mathcal{P}_{\alpha}\}{}_{\alpha=1}^{2^{2n}}$ to obtain the average value. Nevertheless, for the sake of simplicity of the derivations we will maintain the use of $2^{2n}$.
@@
To gain a better grasp of how RC operates, let us express $\mathcal{K}_{\text{cliff}}$ as $\mathcal{K}_{\text{cliff}}=\mathcal{U}_{\text{cliff}}\mathcal{N}$, where $\mathcal{N}$ represents the error channel that encompasses all errors. This decomposition always exists when the dynamics of the gate constitute a completely positive map. Substituting it into the twirling formula, we obtain:
$K_{\text{cliff}}^{RC}$ $=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}^{\prime}\mathcal{U}_{\text{cliff}}\mathcal{N}\mathcal{P}_{\alpha}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}^{\prime}\mathcal{U}_{\text{cliff}}\mathcal{P}_{\alpha}\mathcal{P}_{\alpha}\mathcal{N}\mathcal{P}_{\alpha}$
$=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{U}_{\text{cliff}}\mathcal{P}_{\alpha}\mathcal{N}\mathcal{P}_{\alpha}=\mathcal{U}_{\text{cliff}}\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{N}\mathcal{P}_{\alpha}.$ (5)
@@
Hence, twirling a noisy Clifford gate with $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\alpha}^{\prime}$ is equivalent to twirling its noise channel $\mathcal{N}$ with the same $\mathcal{P}_{\alpha}$ on both sides ($\mathcal{P}_{\alpha}^{\prime}=\mathcal{P}_{\alpha}$). Furthermore, we can utilize the fact that Pauli operators form a complete basis, allowing to express any $\mathcal{N}$ as $\mathcal{N}=\sum_{\alpha,\beta}\mathcal{N}_{\alpha,\beta}P_{\alpha}\otimes P_{\beta}^{*}$ It can be shown that, after Pauli twirling, $\mathcal{N}$ transforms into
$\mathcal{N}^{RC}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{N}\mathcal{P}_{\alpha}=\sum_{\alpha}\mathcal{N}_{\alpha,\alpha}\mathcal{P}_{\alpha}.$ (6)
@@
In this transformation, only the diagonal elements remain. This resulting channel is referred to as a Pauli error channel, and it corresponds to randomly applying a Pauli unitary $\mathcal{P}_{\alpha}$ with a probability of $\mathcal{N}_{\alpha,\alpha}$. Thus, regardless of the original structure of $\mathcal{N}$, Pauli twirling turns it into a Pauli channel. As we show later, this is not the case in PST.
@@
Figure 2: (a) Randomized compiling (RC, AKA Pauli twirling) replaces the original Clifford (e.g. a CNOT) with a sum of Pauli-twirled Clifford gates. RC can only mitigate coherent errors in Clifford gates. (b) Our proposed Pseudo Twirling (PST) scheme can mitigate uncontrolled coherent errors, such as crosstalk, in a general gate. This is achieved by modifying the signs of some of the pulses in the driving field $H_{d}(t)$ that generates the gate according to the rule $H_{d}(t)\to P_{\alpha}H_{d}(t)P_{\alpha}$. (c) Because PST is not restricted to Clifford gates, it can be implemented at various levels of a circuit: at the entire circuit level (orange lines), gate-level (green lines), and intra-gate level (purple lines).
II.2 Introducing the Pseudo Twirling Protocol